You purchased a machine for $1.07 million three years ago and have been applying​ straight-line depreciation to zero for a​ seven-year life. Your tax rate is 40%.

If you sell the machine today​ (after three years of​ depreciation) for $700,000​, what is your incremental cash flow from selling the​ machine?
Answer #1

Depreciation for three years = (Cost of asset * three years)/ Life of asset

Depreciation for three years = (1,070,000*3)/7

Depreciation for three years = 458,571.4286

-------------------------------------

Book value = 1,070,000 - 458,571.4286

Book value = 611,428.5714

----------------------------------------------

Tax on profit on sale = (700,000 - 611,428.5714)*40%

Tax on profit on sale = 35,428.57

----------------------------------------------------------------

Incremental cash flow = 700,000 - 35,428.57

Incremental cash flow = 664,571.4286